# Edit2Restore: Few-Shot Image Restoration via Parameter-Efficient Adaptation of Pre-trained Editing Models

### v1

@@ 1 Introduction
Image restoration, recovering high-quality images from degraded observations, has traditionally relied on training specialized models for each degradation type, requiring thousands of paired training examples per task. Recent all-in-one restoration methods [airnet2022, promptir2023, instructir2024] attempt to handle multiple degradations within a unified framework, yet still demand extensive paired datasets and specialized architectures trained from scratch. For example, InstructIR [instructir2024] requires training a restoration network with over 10,000 synthetically generated text prompts and thousands of image pairs per degradation to achieve state-of-the-art results.
We pose a fundamental question: Can powerful pre-trained image editing models be efficiently adapted for image restoration with dramatically fewer training examples? Recent advances in image generation and editing have produced remarkable models like FLUX.1 Kontext [fluxkontext2025] with sophisticated understanding of image structure, quality, and semantics learned from enormous datasets. These models excel at text-conditioned image-to-image translation tasks such as object removal, style transfer, and content modification. However, their potential for inverse problems, specifically image restoration, remains largely unexplored in the few-shot regime.
In this work, we demonstrate that FLUX.1 Kontext [fluxkontext2025], a prominent text-conditioned image editing model based on rectified flow matching, can be efficiently adapted for multiple restoration tasks using remarkably few examples: just 16-128 paired images per task. FLUX.1 Kontext is a 12 billion parameter model designed for image-to-image translation, capable of performing diverse editing operations guided by natural language instructions. Our key insight is that the visual understanding already encoded in such models including knowledge about image quality, degradation patterns, and natural statistics can be efficiently transferred to restoration tasks through parameter-efficient fine-tuning (PEFT), without requiring the model to be trained from scratch.
We fine-tune lightweight LoRA [lora2022] adapters on FLUX.1 Kontext, guided by simple text prompts that specify the restoration operation (’remove the noise from the image’, ’remove the haze from the image’, ’remove the rain from the image’). The model performs image-to-image translation: given a degraded input image and a restoration instruction, it generates the corresponding clean output. Unlike InstructIR which trains specialized restoration networks from scratch, we leverage the pre-trained visual understanding already embedded in large-scale image editing models. A single unified LoRA adapter, conditioned on task-specific text prompts, effectively handles multiple degradations including denoising, deraining, and dehazing, all learned from minimal training data.
@@ 2.1 Task-Specific Image Restoration Models
Image restoration aims to recover high-quality images from degraded observations. Early research developed specialized models for specific degradation types, including denoising[DnCNN], deblurring [GoPro, DeblurGAN], low-light enhancement [DeepRetinex, RetinexFormer], dehazing [FFA-Net], deraining [MSFPN], desnowing [DesnowNet, SnowFormer], and raindrop removal [Attentive-GAN].CNN-based methods like NAFNet[NafNet], MPRNet[MPRNet], and HINet[HINet], followed by transformer based approaches such as Restormer[restormer2022], SwinIR[swinir2021], and Uformer[Uformer] advanced the field through better architectural designs. However, these methods optimize pixel-level metrics, using $\ell_{1}$ or $\ell_{2}$ losses, requiring thousands of paired data per task and fall short in naturalness (perceptual quality).
@@ 2.2 Multi-Task Image Restoration Models
Recent work has explored models capable of handling multiple restoration tasks. Among these, AirNet [airnet2022] employed contrastive learning to distinguish degradation types. Prompt-based methods [promptir2023, instrcut-ipt, prompt_in_prompt, ProRes, instructir2024] introduced learnable parameters or natural language instructions to encode degradation-specific information, guiding restoration backbones like NAFNet [NafNet] and Restormer [restormer2022]. For instance, InstructIR [instructir2024] uses human-written text prompts to achieve strong results across multiple tasks. However, these methods train restoration networks from scratch, requiring thousands of paired images per task and substantial computational costs. Only recent work like LoRA-IR [lorair] has explored parameter-efficient adaptation through low-rank techniques.
With the success of diffusion models in generation, several works leverage diffusion priors for restoration [DA-CLIP, DiffBIR, autodir, mperceiver, StableSR, DiffRestorer]. While effective, most diffusion-based approaches are computationally expensive and require large data.
@@ 2.3 Text-Conditioned Image Editing Models
Recent text-conditioned image editing models excel at modifying images according to natural language instructions while preserving unrelated content. Early work like InstructPix2Pix [brooks2022instructpix2pix] demonstrated instruction-following editing, while recent models achieve state-of-the-art quality: FLUX.1 Kontext [fluxkontext2025] employs a dual-stream architecture processing image and text separately, Qwen-Image-Edit [wu2025qwenimagetechnicalreport] offers a competitive alternative, and unified models [chen2025janus, ma2024janusflow, xiao2024omnigen] handle both understanding and generation within single architectures.
@@ 2.4 Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) methods[zhou2022coop, lora2022, waveft] enable adapting large pre-trained models to new tasks while training only a small fraction of parameters. LoRA (Low-Rank Adaptation) [lora2022] injects trainable low-rank matrices into model layers, achieving strong performance in language and vision tasks with minimal parameters.
@@ 3 Edit2Restore: Proposed Methodology
We propose a PEFT approach for adapting a pre-trained image editing model to text-conditioned multi-task image restoration with minimal training data. Our method leverages FLUX.1 Kontext [fluxkontext2025], a 12 billion parameter flow matching model designed for text-conditioned image-to-image translation, and adapts it to multi-task image restoration through Low-Rank Adaptation (LoRA) [lora2022]. Figure 1 illustrates our overall framework.
@@
The model consists of three main components that work together to generate images. A variational autoencoder with 16 latent channels encodes images into latent representations $\mathbf{z}=\mathcal{E}(\mathbf{x})$ and decodes them back to pixel space $\mathbf{x}=\mathcal{D}(\mathbf{z})$. Two text encoders, CLIP[clip] and T5[T5], process the text prompt $\mathbf{t}$ into embeddings that condition the generation process. Finally, a flow matching transformer with a hybrid architecture combining double-stream and single-stream transformer blocks operates on the latent space, performing iterative refinement guided by text conditioning.
Unlike traditional diffusion models that predict noise, FLUX.1 Kontext uses rectified flow matching [rectifiedflow2023, lipman2023flow] which learns to predict velocity fields. The forward process constructs a linear interpolation path between clean latent $\mathbf{z}_{\text{clean}}$ and Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$:
@@ 4.1 Experimental Setup
Datasets. We construct small training sets by sampling paired images from established restoration benchmarks. For denoising, we sample from BSD400 [bsd400]; for deraining, from Rain100L [rain100l]; and for dehazing, from the Reside (outdoor) SOTS [sots] dataset. For denoising, we ensure balanced coverage across noise levels by generating approximately equal numbers of degraded images at $\sigma=15,25,50$.
@@
Implementation details. All experiments are conducted on a single NVIDIA H100 GPU. The LoRA rank and batch size are set to $64,4$ respectively for all experiments. Each training run is performed for 1,920 iterations across all experimental configurations to ensure fair comparison. For multi-task training with three degradation types (denoising, deraining, dehazing), this corresponds to approximately 640 iterations per task when using 32 images per task (96 total images with batch size 4). We employ the AdamW optimizer [adamw] with learning rate $1\times 10^{-4}$ for the transformer LoRA parameters. When fine-tuning the text encoder, we use a lower learning rate of $5\times 10^{-6}$ to prevent catastrophic forgetting of pre-trained text representations. Weight decay is set to $1\times 10^{-4}$ for transformer parameters and $1\times 10^{-3}$ for text encoder parameters. We apply a constant learning rate schedule with 500 warmup steps. Training is performed in mixed precision (bfloat16) with gradient checkpointing enabled to reduce memory consumption.
@@
Evaluation metrics. We report three complementary metrics to assess restoration quality from different perspectives: (1) FID (Fréchet Inception Distance) [fid] measures distribution-level similarity between restored and ground truth images, capturing overall perceptual quality and realism; (2) CMMD (CLIP Maximum Mean Discrepancy) [cmmd] evaluates feature-space similarity using maximum mean discrepancy, providing a robust measure of distribution matching; (3) CLIP-IQA [clipiqa] assesses perceptual image quality using CLIP-based features, aligning with human perception of visual quality. These metrics prioritize perceptual quality over pixel-perfect reconstruction, consistent with our approach’s focus on leveraging pre-trained generative priors rather than optimizing for PSNR/SSIM.

### v2

@@ 1 Introduction
Image restoration, recovering high-quality images from degraded observations, has traditionally relied on training specialized models for each degradation type, requiring thousands of paired training examples per task. Recent all-in-one restoration methods [21, 33, 12] attempt to handle multiple degradations within a unified framework, yet still demand extensive paired datasets and specialized architectures trained from scratch. For example, InstructIR [12] requires training a restoration network with over 10,000 synthetically generated text prompts and thousands of image pairs per degradation to achieve state-of-the-art results.
We pose a fundamental question: Can powerful pre-trained image editing models be efficiently adapted for image restoration with dramatically fewer training examples? Recent advances in image generation and editing have produced remarkable models like FLUX.1 Kontext [5] with sophisticated understanding of image structure, quality, and semantics learned from enormous datasets. These models excel at text-conditioned image-to-image translation tasks such as object removal, style transfer, and content modification. However, their potential for inverse problems, specifically image restoration, remains largely unexplored in the few-shot regime.
In this work, we demonstrate that FLUX.1 Kontext [5], a prominent text-conditioned image editing model based on rectified flow matching, can be efficiently adapted for multiple restoration tasks using remarkably few examples: just 16-128 paired images per task. FLUX.1 Kontext is a 12 billion parameter model designed for image-to-image translation, capable of performing diverse editing operations guided by natural language instructions. Our key insight is that the visual understanding already encoded in such models including knowledge about image quality, degradation patterns, and natural statistics can be efficiently transferred to restoration tasks through parameter-efficient fine-tuning (PEFT), without requiring the model to be trained from scratch.
We fine-tune lightweight LoRA [15] adapters on FLUX.1 Kontext, guided by simple text prompts that specify the restoration operation (’remove the noise from the image’, ’remove the haze from the image’, ’remove the rain from the image’). The model performs image-to-image translation: given a degraded input image and a restoration instruction, it generates the corresponding clean output. Unlike InstructIR which trains specialized restoration networks from scratch, we leverage the pre-trained visual understanding already embedded in large-scale image editing models. A single unified LoRA adapter, conditioned on task-specific text prompts, effectively handles multiple degradations including denoising, deraining, and dehazing, all learned from minimal training data.
@@ 2.1 Task-Specific Image Restoration Models
Image restoration aims to recover high-quality images from degraded observations. Early research developed specialized models for specific degradation types, including denoising[47], deblurring [32, 19], low-light enhancement [42, 7], dehazing [35], deraining [17], desnowing [27, 10], and raindrop removal [34].CNN-based methods like NAFNet[8], MPRNet[45], and HINet[9], followed by transformer based approaches such as Restormer[46], SwinIR[23], and Uformer[41] advanced the field through better architectural designs. However, these methods optimize pixel-level metrics, using $\ell_{1}$ or $\ell_{2}$ losses, requiring thousands of paired data per task and fall short in naturalness (perceptual quality).
@@ 2.2 Multi-Task Image Restoration Models
Recent work has explored models capable of handling multiple restoration tasks. Among these, AirNet [21] employed contrastive learning to distinguish degradation types. Prompt-based methods [33, 38, 22, 30, 12] introduced learnable parameters or natural language instructions to encode degradation-specific information, guiding restoration backbones like NAFNet [8] and Restormer [46]. For instance, InstructIR [12] uses human-written text prompts to achieve strong results across multiple tasks. However, these methods train restoration networks from scratch, requiring thousands of paired images per task and substantial computational costs. Only recent work like LoRA-IR [1] has explored parameter-efficient adaptation through low-rank techniques.
With the success of diffusion models in generation, several works leverage diffusion priors for restoration [29, 24, 18, 2, 40, 48]. While effective, most diffusion-based approaches are computationally expensive and require large data.
@@ 2.3 Text-Conditioned Image Editing Models
Recent text-conditioned image editing models excel at modifying images according to natural language instructions while preserving unrelated content. Early work like InstructPix2Pix [6] demonstrated instruction-following editing, while recent models achieve state-of-the-art quality: FLUX.1 Kontext [5] employs a dual-stream architecture processing image and text separately, Qwen-Image-Edit [43] offers a competitive alternative, and unified models [11, 31, 44] handle both understanding and generation within single architectures.
@@ 2.4 Parameter-Efficient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) methods[49, 15, 4] enable adapting large pre-trained models to new tasks while training only a small fraction of parameters. LoRA (Low-Rank Adaptation) [15] injects trainable low-rank matrices into model layers, achieving strong performance in language and vision tasks with minimal parameters.
@@ 3 Edit2Restore: Proposed Methodology
We propose a PEFT approach for adapting a pre-trained image editing model to text-conditioned multi-task image restoration with minimal training data. Our method leverages FLUX.1 Kontext [5], a 12 billion parameter flow matching model designed for text-conditioned image-to-image translation, and adapts it to multi-task image restoration through Low-Rank Adaptation (LoRA) [15]. Figure 1 illustrates our overall framework.
@@
The model consists of three main components that work together to generate images. A variational autoencoder with 16 latent channels encodes images into latent representations $\mathbf{z}=\mathcal{E}(\mathbf{x})$ and decodes them back to pixel space $\mathbf{x}=\mathcal{D}(\mathbf{z})$. Two text encoders, CLIP[36] and T5[37], process the text prompt $\mathbf{t}$ into embeddings that condition the generation process. Finally, a flow matching transformer with a hybrid architecture combining double-stream and single-stream transformer blocks operates on the latent space, performing iterative refinement guided by text conditioning.
Unlike traditional diffusion models that predict noise, FLUX.1 Kontext uses rectified flow matching [26, 25] which learns to predict velocity fields. The forward process constructs a linear interpolation path between clean latent $\mathbf{z}_{\text{clean}}$ and Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$:
@@ 4.1 Experimental Setup
Datasets. We construct small training sets by sampling paired images from established restoration benchmarks. For denoising, we sample from BSD400 [3]; for deraining, from Rain100L [13]; and for dehazing, from the Reside (outdoor) SOTS [20] dataset. For denoising, we ensure balanced coverage across noise levels by generating approximately equal numbers of degraded images at $\sigma=15,25,50$.
@@
Implementation details. All experiments are conducted on a single NVIDIA H100 GPU. The LoRA rank and batch size are set to $64,4$ respectively for all experiments. Each training run is performed for 1,920 iterations across all experimental configurations to ensure fair comparison. For multi-task training with three degradation types (denoising, deraining, dehazing), this corresponds to approximately 640 iterations per task when using 32 images per task (96 total images with batch size 4). We employ the AdamW optimizer [28] with learning rate $1\times 10^{-4}$ for the transformer LoRA parameters. When fine-tuning the text encoder, we use a lower learning rate of $5\times 10^{-6}$ to prevent catastrophic forgetting of pre-trained text representations. Weight decay is set to $1\times 10^{-4}$ for transformer parameters and $1\times 10^{-3}$ for text encoder parameters. We apply a constant learning rate schedule with 500 warmup steps. Training is performed in mixed precision (bfloat16) with gradient checkpointing enabled to reduce memory consumption.
@@
Evaluation metrics. We report three complementary metrics to assess restoration quality from different perspectives: (1) FID (Fréchet Inception Distance) [14] measures distribution-level similarity between restored and ground truth images, capturing overall perceptual quality and realism; (2) CMMD (CLIP Maximum Mean Discrepancy) [16] evaluates feature-space similarity using maximum mean discrepancy, providing a robust measure of distribution matching; (3) CLIP-IQA [39] assesses perceptual image quality using CLIP-based features, aligning with human perception of visual quality. These metrics prioritize perceptual quality over pixel-perfect reconstruction, consistent with our approach’s focus on leveraging pre-trained generative priors rather than optimizing for PSNR/SSIM.